CH_{3}OH SUB-DOPPLER SPECTROSCOPY

GERMAN YU. GOLUBIATNIKOV, SERGEY P. BELOV and \underline{ALEXANDER V. LAPINOV}, Institute of Applied Physics of RAS, 46 Ulyanov str., 603950 Nizhny Novgorod, Russia.

The methanol torsion-rotation spectrum in the first three torsional states has been measured and analysed for a search of m_{e}/m_{p} variations from comparison of radio astronomical and laboratory frequencies and for studies of systematic velocity motions in star-forming regions. The investigation is based on Lamb-dip measurements with sub-Doppler spectrometer developed at IAP RAS. CH_{3}OH-A and -E frequencies have been obtained with an accuracy of ∼1 kHz at 48−510 GHz for more than 500 transitions in v_{t} = 0, more than 200 transitions in v_{t} = 1 and 100 transitions in v_{t} = 2. For many b-type transitions the removed degeneracy of the CH_{3}OH levels due to different nuclear spin statistics was measured in a form of doublets with unresolved hf structure. Some a-type transitions show resolved spin-rotational splitting.